Status Report

Hubble Space Telescope Daily Report #3065 – 27 Feb 2002

By SpaceRef Editor
February 27, 2002
Filed under , ,



PERIOD COVERED: 0000Z (UTC) 02/26/02 – 0000Z (UTC) 02/27/02

Daily Status Report as of 058/0000Z


1.1 Completed Two Sets of STIS/CCD 8901 (Dark Monitor-Part 1)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the darks. There was no anomalous activity.

1.2 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3)

The WF/PC-2 was used obtain three dark frames every day to provide
data for monitoring and characterizing the evolution of hot pixels. No
problems were encountered.

1.3 Completed STIS/CCD 9131 (Imaging the Host Galaxies of High Redshift
Type Ia Supernovae)

The Space Telescope Imaging Spectrograph (CCD) was used to complete
the snapshot survey of distant galaxies of known redshift which hosted
supernovae {SNe} of Type Ia found via the Supernova Cosmology Project
{SCP}. No problems were seen.

1.4 Completed Two Sets of STIS/CCD 9317 (Pure Parallel Imaging Program:
Cycle 10)

The Space Telescope Imaging Spectrograph (CCD) was used to perform
the default archival pure parallel program for STIS during cycle 10. There
were no reported problems.

1.5 Completed STIS/CCD 8903 (Bias Monitor – Part 1)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at
gain = 4 to build up high-S/N superbiases and track the evolution of hot
columns. There were no problems.

1.6 Completed STIS/CCD 9139 (Variability in the UV Spectrum of 3C 279:
Testing Models for the Gamma-Ray Emission in Blazars)

The Space Telescope Imaging Spectrograph (CCD) was used to
determine whether gamma- rays are produced on the scale of the broad line
region, using multi-epoch UV snapshots of 3C 279 to find and characterize
its LyAlpha variability. This will validate the mirror model for
generating the intense gamma-ray emission from blazars. 3C 279 is the
best-studied blazar from radio to gamma-ray wavelengths, showing frequent
large flares over days to months. Variability in the UV continuum and
LyAlpha will directly reveal any coupling between the jet ionizing flux and
the broad-line region emission, providing clues to the physics and
energetics of all radio-loud AGN. All observations completed without
reported incident.

1.7 Completed Five Sets of WF/PC-2 9318 (POMS Test Proposal: WFII
Parallel Archive Proposal Continuation)

The WF/PC-2 was used to perform the generic target version of the
WFPC2 Archival Pure Parallel program. The program was used to take
parallel images of random areas of the sky, following the recommendations
of the Parallels Working Group. There were no reported problems.

1.8 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor)

The Space Telescope Imaging Spectrograph (MA2) was used to perform
daily monitoring of the NUV MAMA detector dark noise in order to monitor
the effects of thermal changes on the NUV dark rate. No problems occurred.

1.9 Completed STIS/CCD/MA1 9182 (The Thermal Glow of Gamma Ray Pulsars)

The Space Telescope Imaging Spectrograph (CCD and MA1) was used to
perform time- and energy- resolved UV observations of two young Gamma-ray
pulsars. The targets, the Vela and Geminga pulsars, are the brightest
neutron stars spanning the critical 10^4-10^6y age range, where the thermal
surface emission is detectable. The UV lightcurves obtained will provide
phase-resolved colors that will isolate this thermal emission and test the
equation of state and fundamental interactions of matter at supernuclear
densities. In addition, combining the UV results with IR/optical and X-ray
data, the proposers will map spectral variations in the non-thermal
pulsations and probe the acceleration physics of these GeV gamma-ray
machines. All observations completed without incident.

1.10 Completed STIS/MA1 8920 (Cycle 10 MAMA Dark Measurements)

The Space Telescope Imaging Spectrograph (MA1) was used to perform
the routine monitoring of the MAMA detector dark noise, and is the primary
means of checking on health of the MAMA detectors systems through frequent
monitoring of the background count rate. The proposal completed with no
reported anomalous activity.

1.11 Completed STIS/MA1/MA2 8918 (MAMA Full Field Sensitivity Monitor &
PSF Check)

The Space Telescope Imaging Spectrograph (MA1 and MA2) was used
monitor the sensitivity of the MAMA detectors over the full field. This
will be achieved by observing globular cluster NGC6681 twice during Cycle
10 and then compared with similar data obtained in Cycles 7, 8 and
9. There were no reported problems.

1.12 Completed WF/PC-2 9580 (Hubble Heritage Observations of NGC 4319)

The WF/PC-2 was used to observe NGC 4319 for a Hubble Heritage Team
release – data will be released as a Public Relations image. No problems
were reported.

1.13 Completed STIS/CCD 9066 (Closing in on the Hydrogen Reionization
Edge of the Universe)

The Space Telescope Imaging Spectrograph (CCD) was used in parallel
constrain the Hydrogen reionization edge in emission that marks the
transition from a neutral to a fully ionized IGM at a predicted
redshifts. The proposal completed uneventfully.

1.14 Completed Two Sets of STIS/CCD 9088 (Next Generation Spectral
Library of Stars)

The Space Telescope Imaging Spectrograph (CCD) was used to produce
a “Next Generation” Spectral Library of 600 stars for use in modeling the
integrated light of galaxies and clusters by using the low dispersion UV
and optical gratings of STIS. The library will be roughly equally divided
among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling
the entire HR-diagram in each bin. Such a library will surpass all extant
compilations and have lasting archival value, well into the Next Generation
Space Telescope era. No problems occurred.


2.1 Guide Star Acquisitions:

Scheduled Acquisitions: 11

Successful: 11

Scheduled Re-acquisitions: 5

Successful: 5

2.2 FHST Updates:

Scheduled: 28

Successful: 28

2.3 Operations Notes:

Using ROP SR-1A, the SSR-1A EDAC error counter was cleared four

The Voice Distribution Boxes in the STOCC area were down from
057/1025Z until 057/1217Z.

SSA transmitter #2 was turned on and off for the interval
057/1030Z to 057/1039Z per ROP IC-2.

The CCS engineering status buffer limit for SESBSLD (=12 when the
HGA torque was exceeded) was modified at 057/1821Z, using ROP DF-18A.

A TTR was written for a required re-transmit during an SI C&DH load
at 057/204739Z. Accordingly, SI C&DH errors were reset at 057/2053Z as
directed by ROP NS-05.


Servicing Mission 3B, is currently scheduled for 6:48 a.m. on
February 28. The countdown continues on schedule, although the weather is
forecasted to be just above the limit at launch time.

Continuation of normal science observations and calibrations.

SpaceRef staff editor.